RC Meeting and Run Plan, Thursday, February 20th 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

Thursday 20th Feb. 2020.

General updates

  • Pretty smooth running since yesterday. Yesterday afternoon the beam was very stable, becoming more trippy over time but not worse than usual. Overall we've had very productive shifts.
  • The only hiccup was with the target around midnight last night, about an hour after the most recent NMR had been taken (55.3%). The correction coil had gone off, and after correcting that problem 2 more NMRs were apparently taken which dropped the polarization to 52%. So after talking about it this morning, we decided that the lesson learned for the future is that if you're concerned about some issue and think you need to take more than one NMR every 5 h, please call the target expert on call before doing so. In last night's case, the correction coil should not have affected the polarization so it wasn't necessary to take more NMRs.

Pending problems

  • ESR compressor crash at 13:51. All spectrometer magnets went down. Helium levels empty, LN2 levels OK. Waiting for magnets to fill... At least it happened while our hall was open fixing the ice problem anyway- Murphy slipped up this time!
  • Still need 2 shifts filled: one on Saturday day and one on Saturday swing. Please sign up for them!!
  • HMS Q1 ice line (see https://logbooks.jlab.org/entry/3788095).
    • Controlled access today at 13:45 (radcon scheduled for 13:15). Techs fixed the stuck-open LN2 return valve (see https://logbooks.jlab.org/entry/3788243), installed tarps to divert meltwater. Kepco ps came out after radcon survey.
  • Offline task of reference cell pressure calibration needs to get done. Brad & Todd are on it -- Done https://logbooks.jlab.org/entry/3788269
  • Display orbit lock status on live display screen (Brad) -- Ops SW folks contacted
  • Ion chamber trip thresholds manipulation by MCC: https://logbooks.jlab.org/entry/3785352 https://logbooks.jlab.org/entry/3783965 https://logbooks.jlab.org/entry/3784607
  • Raster energy alarm, possible cause: it's initialized for 10.43 GeV not 10.383 GeV.
    • The working hypothesis is that the mismatch between the raster GUI energy (10.43 GeV) and the Tiefenbach energy (10.386 GeV) is the reason for our nuisance alarms. I told him I was skeptical of this explanation, and thought it was probably transients, but what the hell, why not change the raster GUI energy the 1/2% to match the actual beam energy anyway. I had to go back to the counting room to give him the process variable name (ecEHCFRHALLC:p). But I noticed that the alarm limits on this are 10 times tighter than the 0.044 GeV actual difference between 10.43 and 10.386 GeV. So it should alarm all the time if that's the problem. And then we had a raster energy alarm while we were looking at the GUI, and the reason was that this PV was "inf", ie it was a readout transient, not a real issue. The crew chief said he'd talk to software about changing the raster energy anyway, and look also for a less noisy PV to use for energy.
    • Brad made some improvements, see https://logbooks.jlab.org/entry/3788240
  • Harp travel problem can't be fixed without taking the harps down (out). Limit switch is already at end of clearance. Way bigger job than we thought (so punt).

Not sure if it is a problem

  • HMS Dipole interlock during Saturday owl shift: it's not known if a specific problem on dipole triggered the interlock but reset was successful.

Fixed problems

  • Low longitudinal polarization seems to be fixed!
* 24 hour period oscillations in raster current (6A out of 25A). No problem, see Mark Jones' HCLOG https://logbooks.jlab.org/entry/3788235 where he checked, found the raster size is constant. Take a breath of relief!

Opportunistic Problems to get done at some point

  • E calibration (day shift, no target, tune beam)
  • BCM calibration (to 40 uA, target out, day shift)
  • Note there may need to be a recesiation of the photocathode in 10-14 d which could take 10-12 h, according to what Joe Grames said at Wednesday's 1:30 meeting.
    • The injector group may just move the spot- it remains to be seen.
    • This could be checked when we go down for 2-3 h in the next few days to fix a PSS problem with the lasers in the injector

Opportunistic Access Jobs

  • Walter would like to pump out some of the insulating vacuum lines (needs ~4-5 h). Note that the SHMS xfer line IV was done today during the ESR crash- still need to do the HMS.


Recent polarization measurements

Current Run Plan

  • Taking production data with 3He. Stay in longitudinal until further notice. NMR every 5 hrs.
  • Next Moller measurement next week sometime.
  • If beam goes away for more than 2h, or if during the downtime, MCC does operations that may change then beam tuning (except work on transport magnets), then do beam recovery procedure (do YAG viewer part only if you are suspicious of the harp scan results)
  • Check the position locks are on.
  • Check CODA is running normally (had one strange hangup lately)
  • CHeck online replays shortly after starting runs, & full replay results too.
  • Make sure all the beam specs are followed (e.g. no raster size below 4mm on 3He cell).
  • Always take 120Hz of EDTM
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.


!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Future

  • We will return to the 3.4/3.5 GeV spectrometer momentum settings (angles stay the same) sometime this weekend probably, to get more stats there once we get enough at the higher rate setting we're at now.
  • The transition to d2n will probably happen sometime the week of March 9-13.
  • For the time being we will run exclusively with longitudinal polarization. General idea: if target polarization is about equal for transverse vs. longitudinal, alternate 4 hours of transverse and 20 hours of longitudinal data taking. Do NMR before and after transverse, and every 5 hours during longitudinal.
  • Production at 30 degrees. We may want to start from HMS/SHMS=-2.9/-2.6 momentum setting, complete this and then (maybe) eventually go back to HMS/SHMS=-3.5/-3.4 momentum.

When there is a down:

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If SHMS dipole trips again, take a screenshot of the PSU window before resetting the interlock and post it to hclog.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[1]
  • Temporary Laser instructions can be found here [2]
  • Do not change SHMS momentum until we fixed the current lead valve

Planned Accelerator Configuration Changes

  • Note there will likely be a recesiation of the photocathode in 10-14 d which could take 10-12 h, according to what Joe Grames reported at the 2/19 1:30 meeting.
    • The injector group may just move the spot- it remains to be seen.
    • This could be checked when we go down for 2-3 h in the next few days to fix a PSS problem with the lasers in the injector
  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links